### v1

# LightPAL: Lightweight Passage Retrieval for Open Domain Multi-Document Summarization

Abstract
Open-Domain Multi-Document Summarization (ODMDS) is crucial for addressing diverse information needs, which aims to generate a summary as answer to user’s query, synthesizing relevant content from multiple documents in a large collection. Existing approaches that first find relevant passages and then generate a summary using a language model are inadequate for ODMDS. This is because open-ended queries often require additional context for the retrieved passages to cover the topic comprehensively, making it challenging to retrieve all relevant passages initially. While iterative retrieval methods have been explored for multi-hop question answering (MQA), they are impractical for ODMDS due to high latency from repeated large language model (LLM) inference for reasoning. To address this issue, we propose LightPAL, a lightweight passage retrieval method for ODMDS that constructs a graph representing passage relationships using an LLM during indexing and employs random walk instead of iterative reasoning and retrieval at inference time. Experiments on ODMDS benchmarks show that LightPAL outperforms baseline retrievers in summary quality while being significantly more efficient than an iterative MQA approach.
LightPAL: Lightweight Passage Retrieval for Open Domain Multi-Document Summarization
Masafumi Enomoto1, Kunihiro Takeoka1, Kosuke Akimoto1, Kiril Gashteovski2, Masafumi Oyamada1 1NEC Data Science Research Laboratories, 2NEC Laboratories Europe Correspondence: masafumi-enomoto@nec.com
1 Introduction
footnotetext: This paper is under review.
Figure 1: Overview of LightPAL, a lightweight retrieval method for Open-domain Multi-Document Summarization (ODMDS). To retrieve passages required for abstract information needs, LightPAL performs the following processes: 1) Passage Graph Construction: Evaluate the relevance between passages using LLM generation probabilities and create edges between passages offline. 2) Retrieval: After retrieving initial passages using an off-the-shelf retriever, perform a Random Walk on the Passage Graph to retrieve informative context passages that are referenced by many passages. LightPAL can lightly retrieve passages that are the context for already retrieved passages without iterative reasoning and retrieval, while existing off-the-shelf retrievers cannot do so without iteration.
Summarization is a task that aims to generate text that preserves the most informative content from input documents. As an advanced version of the summarization task, Open-Domain Multi-Document Summarization (ODMDS) Giorgi et al. (2023) has recently been introduced, where the inputs are a large collection of documents, and the summary should be an answer to a query that reflects the information needs of the user. For example, given a query such as “What is the story of THE RECRUIT, an SF novel?”, a summarization system needs to synthesize information from multiple sections of the novel, which may be scattered across different documents within a large collection. In this way, ODMDS requires: (1) generating a summary in response to queries that ask for an overview or open-ended questions, i.e., queries that demand a summary as an answer, and (2) synthesizing the content of multiple documents related to the query within the large document collection. Users typically have diverse information needs, making it impractical to precompile a document collection that caters to all possible queries. Consequently, the ODMDS task plays a vital role in information-seeking scenarios where the system must adapt to the specific requirements of each user.
A practical approach to ODMDS is the retrieve-then-summarize framework Giorgi et al. (2023), where a retrieval model first finds relevant passages for the query, and then a language model generates a summary using the query and retrieved passages. Producing a good summary crucially depends on retrieving comprehensive and relevant information in the first step, as low coverage or irrelevant information leads to poor summaries. Existing work Zhou et al. (2023) has employed sparse and pre-trained dense retrievers, similar to open-domain question answering Zhu et al. (2021). However, these retrievers are inadequate for the ODMDS because the information needs in ODMDS are abstract since they involve overviews or open-ended questions. When the information need is abstract, the query is not expressed as a concrete question like a factoid. Consider again the query, “What is the story of THE RECRUIT, a SF novel?” (see Figure 1). This query entails potential information needs such as “Who are the characters?”, but without reading the story, it is impossible to know the entities needed to concretize such questions to search relevant passages. Therefore, existing retrievers search for passages containing the novel title “THE RECRUIT,” but they cannot search for information that addresses potential information needs such as “Who is Wayne in THE RECRUIT?” which arises after reading the passage. To answer such abstract information needs using existing retrievers, recursive search is required, where initially retrieved information is used for reasoning to search for further information.
Iterative retrieval and reasoning methods have been primarily proposed for multi-hop question answering (MQA), with recent methods leveraging the reasoning capability of large language models (LLMs). For example, Trivedi et al. (2023) let LLM generate Chain-of-Thought Wei et al. (2022) to make queries to retrieve new documents, while Khalifa et al. (2023) traverse the graph of hyperlinked Wikipedia paragraphs along the path that maximized the generation probability of the given query. In this study, we applied these approaches to ODMDS for the first time and confirmed that both retrieval performance and summary quality improve in the ODMDS task (see Section 5). However, these methods are impractical for ODMDS because the number of passages relevant to the query is larger than that of MQA. This is due to the fact that ODMDS requires diverse information to answer overview queries. For instance, in the Story dataset for ODMDS benchmark Zhou et al. (2023), the average number of relevant passages is 9, while in the HotPotQA Yang et al. (2018) dataset for Multi-hop QA, at most two passages are required for an answer. Consequently, numerous iterations are necessary to gather the relevant passages for the ODMDS task, and performing LLM inference at each iteration leads to high latency. Therefore, a more efficient approach is needed to balance summary quality and computational cost for ODMDS.
In this paper, we propose LightPAL, a lightweight passage retrieval method for ODMDS. It constructs a graph representing passage relationships using a LLM during indexing and then employs random walk on this graph via Personalized PageRank (PPR) Jeh and Widom (2003a) instead of iterative reasoning and retrieval at inference time. During indexing, the LLM evaluates if a passage could serve as the context for another, creating links between passages with high generation probabilities. The generation probabilities of the LLM can capture relevance between passages by considering their diverse styles and domains in the large passage collection. At retrieval, an initial set of relevant passages is retrieved using an arbitrary model. PPR then calculates the probability of reaching other passages from the initial set via a random walk on the graph. Highly probable passages are retrieved as additional context, enabling low-latency retrieval of diverse, relevant information without LLM inference during runtime.
We evaluated LightPAL on two benchmarks: ODSum Zhou et al. (2023) for ODMDS and Querysum Liu et al. (2024) for query-focused multi-document summarization (qMDS), which we adapted for ODMDS by sampling irrelevant passages from the same dataset to create a large passage collection, as qMDS assume only query-relevant passages. In the retrieve-then-summarize approach, we evaluated the quality of generated summaries in terms of relevance to reference summaries, diversity, and comprehensiveness using LLMs. Since no conventional retrieval methods specifically designed for the ODMDS task, we compared LightPAL with a pre-trained dense retriever and confirmed that our proposed method outperforms it as an unsupervised method. Additionally, we compared LightPAL to PromptRank Khalifa et al. (2023), an unsupervised iterative reranking method for MQA. Our method demonstrated favorable summary performance while achieving significantly lower latency, about 1k times faster than PromptRank  [^1] Mistral-7B-v0.1 is used as an LLM in PromptRank..
2 Related Works
2.1 Open-domain Multi-document Summarization
Open-Domain Multi-Document Summarization (ODMDSGiorgi et al. (2023); Zhou et al. (2023) is a task that involves extracting and summarizing diverse information from multiple documents into a concise and informative summary. It generalizes the query-focused summarization (QFS) and multi-document summarization (MDS) tasks. Traditionally, MDS task Ma et al. (2023), with applications like summarizing news articles Fabbri et al. (2019), scientific literature reviews Lu et al. (2020), and product reviews Wang and Ling (2016); Ganesan et al. (2010), focused on creating a single summary from a predefined set of documents. To cater to different user information needs, QFS Kulkarni et al. (2020); Liu et al. (2024), was introduced, where a user query is added to generate summaries relevant to the query.
Building upon this, Giorgi et al. (2023) introduced ODMDS as a task that involves retrieving relevant documents from open-domain information sources and summarizing them to answer a given query. This research created a pseudo-ODMDS dataset using existing MDS data to evaluate the capabilities of a retrieve-then-summarize pipeline. The results revealed a decline in summary quality due to the retrieval phase. More recently, Zhou et al. (2023) created a benchmark for the ODMDS task and evaluated the performance of LLMs as summarization models. Although this work has focused on the impact of different retrieval models on summary quality, there has been no work that enables structured retrieval by associating passages at the indexing phase to address queries seeking an overview.
2.2 Retrieval for Multi-hop/document QA
Several approaches have been developed for Multi-hop Question Answering (MQA) tasks Mavi et al. (2022), which involve gathering and integrating information from multiple sources to obtain an answer. These methods typically involve iterative reasoning and retrieval. Das et al. (2019) utilizes the internal state of the reader model to update the query vector and rerank passages. Qi et al. (2019) uses the reader to extract text spans from the retrieved passages, and then uses these text spans as queries to retrieve additional passages.
Recently, retrieval using an LLM iteratively is commonly used for MQA tasks. Khalifa et al. (2023) proposes an unsupervised re-ranking method for MQA that focuses on exploring relevant hyperlinked passages and assigning them higher scores based on the probabilities of generating the given query by LLMs. Trivedi et al. (2023) utilizes the Chain-of-Thoughts technique for retrieving potentially related passages, and Wang et al. (2024) builds a passage graph from structured documents and traverses the graph to retrieve them by using an LLM agent iteratively. The iterative methods are impractical for ODMDS regarding latency, as mentioned in Section 5.3.
2.3 Retrieval for Long-form QA
Long-form Question Answering (LFQAFan et al. (2019) is similar to ODMDS, requiring paragraph-length responses to complex, open-ended questions. In LFQA, multiple pieces of evidence must be retrieved from a large corpus to answer the query. Existing research divides retrieval systems into non-interactive and interactive methods. Non-interactive methods Su et al. (2022); Krishna et al. (2021) use off-the-shelf dense/sparse retrievers to search the corpus with the initial query. Interactive methods Nakano et al. (2021); Qin et al. (2023) learn an agent that performs interactive web searches like humans. Recent web agent research by Qin et al. (2023) aims to deepen topic understanding and create informative answers by issuing follow-up questions using relevant searched information, similar to human behaviors. Our method shares this motivation but avoids their drawbacks: (1) requiring web-browsing behaviors to train the agent and (2) being inapplicable to non-web-based retrieval systems. Our unsupervised approach can be applied to any retrieval model, offering a wider range of applications.
3 ODMDS Task
Task Formulation. Open-Domain Multi-Document Summarization (ODMDS) is formally defined as follows: given a natural language query $q$ and a large collection of passages $D=\{d_{i}\}_{i}$, generate a summary $S$ corresponding to the query $q$.
Approach. Following prior work Zhou et al. (2023), we adopt retrieve-then-summarize approach for the ODMDS task  [^2] While some LLMs, such as Gemini Pro 1.5(https://deepmind.google/technologies/gemini/pro/), can handle up to 1M input tokens, the size of the passage collection is generally unbounded, making retrieval a necessity.. First, we retrieve a set of passages $D_{q}\subset D$ relevant to the query $q$ from the collection $D$. Then, we input the retrieved text and query to a language model $LM_{sum}$ to generate an abstractive summary $S=LM_{sum}(q,D_{q})$.
4 LightPAL: Proposed Method
We propose a retrieval method for addressing open-ended ODMDS queries. Our approach retrieves relevant and diverse passages while avoiding iterative retrieval and language model inference at runtime. As shown in Figure 1, The method consists of two phases:
  1. 1.
    Graph Construction: Pre-constructing a graph representing passage relationships using a large language model (LLM), linking passages with high conditional generation probabilities as potential context for a passage.
  2. 2.
    Retrieval: Retrieving an initial set of relevant passages with a conventional retriever, then employing random walk via Personalized PageRank (PPR) Jeh and Widom (2003a) on the pre-built graph. PPR calculates probabilities of reaching other passages from initial passages, retrieving highly probable ones as additional context without runtime language model inference.
4.1 Passage Graph Construction
The motivation behind our approach is that passages extracted from a larger document may not be self-explanatory and require context from other passages to be correctly understood. Even cohesive passages can benefit from related texts to deepen topic understanding. Naive retrieval systems that only consider query-passage relevance cannot capture these relationships. We aim to address this by creating an graph annotated with passage relationships using the conditional generation probabilities of LLMs. Concretely, for each ordered pair of passages $(d_{i},d_{j})$ in passage set $D$, we calculate the conditional generation probability of passage $d_{j}$ given passage $d_{i}$ as a context score using a language model:
$\text{ContextScore}(d_{i},d_{j})=P_{LM}(d_{j}|d_{i}).$
We construct a passage graph $G$ by creating edges between the passage $d_{i}$ and the top-$k$ passages $\{d_{j}\}_{j}$ with the highest $\text{ContextScore}(d_{i},d_{j})$. Calculating context scores for all passage pairs is computationally prohibitive. Thus, we first retrieve the top-$k^{\prime}$ candidate passages similar to each passage using a lightweight embedding model and then run the language model on these candidates to create edges for the top-$k$ ($k\ll k^{\prime}$) passages with the highest context scores.
We choose the conditional generation probabilities of LLMs to build a passage graph because they can richly determine the relevance between passages by considering the diverse styles and domains of the passages in the passage set $D$. They capture the fluency of the passage sequence, making them suitable for reconstructing relationships between passages originally part of a cohesive document or for building new relationships between passages from different documents that contribute as the context.
4.2 Passage Retrieval
Our proposed method consists of two phases for passage retrieval: 1) Initial Passage Retrieval: Retrieve passages $D_{init}$ directly relevant to the query using any sparse/dense retrieval model. 2) Context Passage Retrieval: Using the initial passages, additionally retrieve context passages $D_{context}$ from the passage graph $G$ that provide contextual information to the initial passages. Finally, we obtain $D_{q}$ as the union of $D_{init}$ and $D_{context}$.
For context passage retrieval, we employ Personalized PageRank (PPR) Jeh and Widom (2003b). PPR calculates the probability of reaching other passages from the initial set $D_{init}$ via a random walk on the graph $G$, highly ranking passages cohesively referenced by many passages. Starting a random walk from the query-relevant initial passages, we expect to explore topics related to the query, and then retrieve highly informative context passages for those topics using PPR. This approach enables low-latency retrieval of relevant information without iterative search using LLM inference during runtime.
5 Experiments
5.1 Settings
Data. To evaluate our approach, we used three datasets: (1) Story dataset from ODSum Zhou et al. (2023), where passages are sections from independent stories, requiring retrieval of cohesive story sections relevant to the query. (2) Meeting dataset from ODSum, where queries are questions about meeting contents. We divided each long meeting transcript (approx. 8,000 tokens) into smaller chunks of around 1,000 characters, treating each chunk as a passage to avoid irrelevant information degrading performance. If meeting transcript is labeled as relevant to query, chunks are also labeled as relevant. We used test set from both Story and Meeting. (3) Querysum Liu et al. (2024), a dataset for query-focused multi-document summarization, consists of general non-factoid questions as queries and Wikipedia paragraphs as associated passages. We selected queries with long reference summaries (70+ words) and multiple relevant passages to test complex questions. Since Querysum does not provide a passage collection $D$ for searching, we created a collection $D$ by randomly extracting 5k Wikipedia paragraphs that are not related to the selected queries and combining them with the associated passages. This ensures that the collection contains a sufficient number of irrelevant passages. To reduce the computational cost of summary generation, we truncated each passage to its first 2,000 characters, assuming that the entity information needed to answer queries is primarily found in the beginning of paragraphs. We refer to the dataset created using the above procedure as QsumOD and use it in our experiments. The statistics of the datasets are shown in Table 1.
Dataset Story Meeting QsumOD
# Queries 260 131 90
# Passages 1190 13074 5803
# PassLen. 3170 898 1391
# SummLen. 1286 1081 675
# Rel. 9.4 170.9 8.9
Rel. % 0.8% 1.3% 0.2%
Table 1: Statistics of the datasets. Abbreviations used: Rel. (Average Relevant Passages), PassLen. (Average Passage Length in characters), SummLen. (Average Reference Summary Length in characters), Rel. % (Relevant Passages Ratio).
Baselines for Passage Retrieval. To investigate the impact of retrieval methods on the final summary quality, we prepared: 1) A pre-trained dense retriever that embeds passages and queries into the same vector space and retrieves the top-K passages most similar to the query. We used the bge-large-en-v1.5  [^3] https://huggingface.co/BAAI/bge-large-en-v1.5 model for embedding as it is a top-tier model in Massive Text Embedding Benchmark  [^4] https://github.com/embeddings-benchmark/mteb 2) PromptRank Khalifa et al. (2023), an unsupervised reranker for Multi-hop QA, which performs iterative reasoning and retrieval by traversing links between passages. PromptRank requires links between passages, so to enable a fair comparison, we employ the passage graph created by our method. Mistral-7B-v0.1  [^5] https://huggingface.co/mistralai/Mistral-7B-v0.1 is used as a LLM in PromptRank. For a fair comparison, both PromptRank and our proposed method initially extracted the top $|D_{init}|$ passages from the $K$ passages retrieved by the dense retriever. Each method, PromptRank and our proposed method, then retrieved the remaining $|D_{context}|=K-|D_{init}|$ context passages using their respective approaches. We set $|D_{init}|$ as top 60% of $K$.
Detailed Settings. In the proposed method, we used the bge-large-en-v1.5 embedding model to obtain the top-100 most similar passages for each passage as candidates. For a fair comparison with PromptRank, we calculated the ContextScore for these passage pairs using Mistral-7B-v0.1, the same model employed by PromptRank. We then constructed a graph by creating edges between the top-5 pairs based on their ContextScore. The ContextScore computation, performed using A100 PCIe 80GB (8 GPUs), took 43 hours for QsumOD, 97 hours for Meeting, and 9 hours for Story datasets. In the Personalized PageRank algorithm, edge weights were uniformly set, and the algorithm was configured to jump uniformly to the top-20 of initial passages $D_{init}$ with a probability of $1-\alpha$, where $\alpha$ was set to 0.2.
Command-r Mixtral
Comp Div Rel Comp Div Rel
Story
PrR 66 67 60 55 60 57
PAL 56 55 55 49 49 57
QsumOD
PrR 51 58 53 53 57 48
PAL 53 57 44 54 54 42
Meeting
PrR 40 47 53 55 60 52
PAL 45 53 39 45 50 53
Table 2: Head-to-head win rates vs. a naive dense retriever for summary quality (Comprehensiveness, Diversity, and Relevance w.r.t reference summary) judged by GPT-4. <50% in red, >50% in blue. Results when the maximum number of retrieved passages are shown. PrR and PAL stand for PromptRank and LightPAL, respectively.
Language Models for Summary Generation. For the ODSUM task, where the number of passages required for generating high-quality summaries is unknown in advance, we used top-tier long-context language models from the RULER benchmark Hsieh et al. (2024) as summary generators to handle a wide range of context lengths: c4ai-command-r-v01  [^6] https://huggingface.co/CohereForAI/c4ai-command-r-v01 and mixtral-8x22B-Instruct-v0.1  [^7] https://huggingface.co/mistralai/Mixtral-8x22B-Instruct-v0.1. For Command-r, we evaluated from 20 to 200 passages for Meeting, from 10 to 100 passages for Story, and from 20 to 100 passages for QsumOD. Due to context limits, Mixtral was evaluated with up to 40 passages for Story and 60 for QsumOD. The prompts used for summary generation are provided in Appendix A.2.
Evaluation Aspects. We conducted a pairwise comparison using an LLM  [^8] We used OpenAI’s gpt-4-turbo-2024-04-09 (https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4) as the evaluator. to investigate whether our proposed method and PromptRank generate higher-quality summaries than the naive approach for the same queries. The LLM greedily decoded and chose between two options (A or B) for the better summary. We evaluated the summaries along the following three dimensions: (1) Relevance: How much important content from the reference summaries is covered. Assessed whether the summary captures the key points and essential information from the reference summaries. (2) Diversity: How varied and insightful the summary is in providing different perspectives on the question. Assessed whether the summary introduces multiple viewpoints and facilitates a broader understanding of the topic. (3) Comprehensiveness: How well the summary covers all aspects of the question. Evaluated whether the summary provides a thorough representation of information that directly addresses the question. The prompts used for evaluation are provided in Appendix A.1.
CPU time GPU time
Data LightPAL LightPAL PromptRank
ST 0.001 s 0.014 s 46.506 s
QS 0.007 s 0.016 s 17.993 s
MT 0.017 s 0.017 s 12.342 s
Table 3: Average latency per query in context passage retrieval in seconds. The CPU used is an AMD EPYC 7302 16-Core Processor (one thread), and the GPU is an A100 PCIe 80GB (8 GPUs). MT, ST, and QS are abbreviations for Meeting, Story, and QsumOD, respectively.
Figure 2: Average win rates by graph construction method (GenProb: Language Model Generation Probability, EmbSim: Passage Embedding Similarity) when the summary generation model is fixed to Command-r. Top: Passage Retrieval w/ LightPAL. Bottom: Passage Retrieval w/ PromptRank.
5.2 Summary Quality Comparison
Table 2 presents the pair-wise comparison of summaries against the naive retriever. Story: The proposed and existing methods outperform the naive dense retriever, with PromptRank having a significant impact (up to 67% win rate for div). LightPAL shows smaller but consistent improvements, except for Mixtral’s Comp/Div, which are mostly tied. QsumOD: Both the proposed and existing methods improve Comprehensiveness and Diversity over the naive dense retriever, especially for Command-r. However, Relevance to reference summaries generally deteriorates, possibly due to the short reference summaries (see Table 1) lacking diverse content relative to the query, resulting in a trade-off between diversity and relevance. Meeting: In general, both the LightPAL and PromptRank result in worse performance compared to the naive dense retriever. This may be attributed to two factors. First, the queries in the Meeting dataset are artificially composed of multiple specific queries in QMSum dataset Zhong et al. (2021). The dense retriever might have already retrieved sufficiently query-relevant passages via such specific queries. Second, the Meeting dataset contains passages in a direct speech style, which may make it more challenging to capture the dependencies between the content. Utterances tend to be more verbose than third-person narrative text, and this verbosity may hinder the understanding of content dependencies. The impact of graph construction methods on performance will be explained in Section 5.4, and the retrieval performance will be discussed in Section 5.5.
In conclusion, PromptRank and LightPAL improve performance in QsumOD and Story datasets, particularly in Comprehensiveness and Diversity. The proposed method, while inferior to PromptRank, still improves performance favorably.
5.3 Latency Comparison in Passage Retrieval
Table 3 shows the average latency for retrieving context passages for a single query. On both CPU and GPU, LightPAL is approximately 1,000 times faster than PromptRank. Also, PromptRank’s latency exceeds 10 seconds, which is prohibitive for interactive RAG systems. The time complexity for computing Personalized PageRank scores is linear in the number of edges in the graph Yang et al. (2024), allowing for practical search time.
Story (top-$k$) QsumOD (top-$k$)
Method 10 20 30 40 50 20 40 60 80 100
Precision@$k$
Naive 29.62 17.54 13.00 10.38 8.67 14.57 9.81 7.50 6.04 5.13
PrR 48.85 32.00 23.49 18.27 14.97 13.35 9.72 8.74 7.61 6.62
PAL 43.15 25.67 17.92 14.29 11.92 13.78 9.31 7.43 6.13 5.14
Recall@$k$
Naive 36.61 42.28 46.81 49.41 51.27 37.05 47.48 52.84 56.40 58.89
PrR 56.52 70.86 77.03 79.88 82.01 33.42 43.36 49.60 52.93 55.27
PAL 50.54 58.45 60.66 64.61 67.39 34.98 44.96 52.95 56.87 59.27
Meeting (top-$k$)
Method 20 40 60 80 100
Precision@$k$
Naive 19.69 16.15 14.39 13.34 12.46
PrR 22.06 19.47 16.82 14.94 13.68
PAL 21.37 18.07 15.89 14.55 13.69
Recall@$k$
Naive 3.02 4.90 6.35 7.67 8.77
PrR 3.49 5.95 7.50 8.77 9.80
PAL 3.28 5.52 7.17 8.56 9.79
Table 4: Comparison of retrieval performance for each method with varying numbers of total passages. PrR and PAL stand for PromptRank and LightPAL, respectively.
5.4 Impact of Graph Construction Method
We analyze the impact of graphs constructed using LLM generation probabilities (GenProb) on summarization quality. As a baseline, we create a graph (EmbSim) that connects each passage to the top-5 passages with the highest cosine similarity based on passage embeddings by bge-large-en-v1.5. Both graphs have the same number of edges per passage.
Figure 2 shows the average win rate of both graphs against the naive retriever in terms of Comprehensiveness, Diversity, and Relevance. Regardless of whether PromptRank or LightPAL is used for context passage retrieval, the GenProb graph outperforms EmbSim on the Story dataset. This is likely because Story passages are sections of a cohesive narrative, and retrieving the entire story is necessary to answer queries. The language model captures the clear sequential relationship between sections of the same story. Furthermore, in PromptRank, the GenProb graph tends to outperform EmbSim on other datasets (Meeting, QsumOD), while this is not the case with LightPAL. These results suggest that the effectiveness of graph construction methods depends on (1) the structure of the passage set and (2) the graph traversal algorithm. At least in PromptRank, which performs iterative reasoning, linking passages using language model generation probabilities (GenProb) positively impacts retrieval.
5.5 Retrieval Performance
To confirm that the improvement in summary quality is due to enhanced retrieval performance, we evaluated Precision@$k$ and Recall@$k$ for the top-$k$ retrieved passages. Precision@$k$ measures the fraction of passages among the top $k$ retrieved results that are actually relevant to the query. Recall@$k$, on the other hand, measures the fraction of all relevant passages that are successfully retrieved within the top $k$ results. Better retrieval performance leads to retrieved passages being more aligned with the reference summary since it was created using query-relevant passages. Consequently, the generated summary from retrieved passages becomes more consistent with the reference summary, which will manifest as an increase in the Relevance metric.
Table 4 presents results when varying the number of top-$k$ passages. Story: LightPAL/PromptRank shows clear improvement over the naive approach, aligning with the improvement in Relevance seen in Section 5.2. QsumOD: LightPAL/PromptRank exhibits little variation from the naive method, which is consistent with the lack of improvement in Relevance. Meeting: LightPAL/PromptRank demonstrates a marginal performance improvement. However, Section 5.2 confirmed that summary quality, especially Relevance, does not improve from the naive approach. This result suggests that our settings to labeling relevant passages for the Meeting dataset may be flawed. Dividing the relevant meeting script into smaller passages and labeling all of them as relevant has likely led to marking many irrelevant passages as relevant, potentially resulting in an inaccurate evaluation of the retrieval performance.
6 Conclusion
In this paper, we proposed LightPAL, a lightweight passage retrieval method for Open-Domain Multi-Document Summarization. Our method pre-constructs a graph representing passage relationships using language models during indexing. At retrieval time, random walk via Personalized PageRank on this graph enables low-latency retrieval of diverse relevant information without runtime LM inference. Experiments on ODMDS and query-focused summarization benchmarks show that our proposed method outperforms baseline retrievers in summary quality while being significantly more efficient than an iterative retrieval and reasoning approach for Multi-hop QA.
7 Limitations
Lack of human evaluation. This paper lacks human evaluation of the generated summaries. Although LLM-based summary evaluations are known to align moderately with human judgments Liu et al. (2023), actual human evaluations are required to further validate the reliability of the results.
Evaluation on datasets other than ODMDS benchmarks. This research has only been evaluated on benchmarks for ODMDS and query-focused multi-document summarization. However, there are other tasks that require long answers based on queries for large document collection. A prominent example is long-form QA Fan et al. (2019), where the queries are complex and open-ended.
Variation of retrieval model (dense/sparse). Our proposed method can use any retrieval model for initial passage retrieval. This paper used a single SOTA dense retriever, i.e., bge-large-en-v1.5. However, the impact of using other pretrained dense retrievers or sparse retrievers like BM25 has not been evaluated.
Variation of LLMs for summary generation. This paper used representative LLMs capable of handling long contexts (c4ai-command-r-v01 and mixtral-8x22B-Instruct-v0.1) and evaluated the quality of summaries when inputting the maximum possible number of passages. However, this methodology increases the inference cost of LLMs. A more practical approach would be to limit the number of retrieved passages, i.e., reduce information recall and increase precision, for summary generation. In such scenarios, models with shorter context lengths may be sufficient for generating summaries. Therefore, a comprehensive evaluation using diverse language models as summarizers is required.
Variation of language model for PromptRank and other iterative approaches. PromptRank is not the only approach that iteratively performs reasoning and retrieval. While PromptRank was chosen as a representative approach for comparison, other methods using Chain-of-Thoughts Trivedi et al. (2023) or LLMs as graph traversal agent Wang et al. (2024) exist. Moreover, there is a trade-off between the capability (size) and inference speed of the language model used for reasoning. Optimizing this trade-off could lead to faster methodologies, but this evaluation has not been performed.
Appendix A Prompts for Generation/Evalaution
A.1 Prompts for pair-wise summary comparison
You will be given a question and two answers (A/B) written for the question.
Your task is to select answer A or B based on their relevance.
Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed.
Your response must be just "A" or "B".
Evaluation Criteria:
Relevance - Selection of important content from the source. The answer should include only important information regarding the question from the source text. Annotators were instructed to penalize answers which contained redundancies and excess information.
Question:
{{Question}}
Source Text:
{{Document}}
Answer A:
{{AnswerA}}
Answer B:
{{AnswerB}}
Evaluation Form ("A"/"B" ONLY):
Please indicate your choice by entering "A" or "B".
Your response:
Listing 1: Prompt for Relevance Evaluation
You will be given a question and two answers (A/B) written for the question.
Your task is to select answer A or B based on their diversity.
Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed.
Your response must be just "A" or "B".
Evaluation Criteria:
Diversity - How varied and insightful is the answer in providing different perspectives on the question? Assess whether the answer introduces multiple viewpoints and facilitates a broader understanding of the topic.
Question:
{{Question}}
Answer A:
{{AnswerA}}
Answer B:
{{AnswerB}}
Evaluation Form ("A"/"B" ONLY):
Please indicate your choice by entering "A" or "B".
Your response:
Listing 2: Prompt for Diversity Evaluation
You will be given a question and two answers (A/B) written for the question.
Your task is to select answer A or B based on their comprehensiveness.
Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed.
Your response must be just "A" or "B".
Evaluation Criteria:
Comprehensiveness - How thorough and detailed is the answer in covering all aspects of the question?
Question:
{{Question}}
Answer A:
{{AnswerA}}
Answer B:
{{AnswerB}}
Evaluation Form ("A"/"B" ONLY):
Please indicate your choice by entering "A" or "B".
Your response:
Listing 3: Prompt for Comprehensiveness Evaluation
A.2 Prompts for summary generation
You are a helpful assistant that gives long answer to question based on long stories. Write an answer based on the following question and the separated sections in the stories. Please write the answer in approximately {{n_summ_words}} words.
# STORY: {{story_idx}}
## SECTION: {passage_idx}}
{{passage}}
... (repeated for all retrieved passages)
QUESTION: {{query}}
SUMMARY:
Listing 4: Prompt for summary generation in Story dataset
You are a helpful assistant that gives long answer to question based on long meetings. Write an answer based on the following question and the separated sections in the meetings. Please write the answer in approximately {{n_summ_words}} words.
# MEETING: {{meeting_idx}}
## SECTION: {passage_idx}}
{{passage}}
... (repeated for all retrieved passages)
QUESTION: {{query}}
SUMMARY:
Listing 5: Prompt for summary generation in Meeting dataset
You are a helpful assistant that gives long answer to question based on documents. Write an answer based on the following question and the documents. Please write the answer in approximately {{n_summ_words}} words.
# DOCUMENT: {{passage_idx}}
{{passage}}
... (repeated for all retrieved passages)
QUESTION: {{query}}
SUMMARY:
Listing 6: Prompt for summary generation in QsumOD dataset
Appendix B Implementation Details
B.1 Used packages
We used rustworkx  [^9] https://github.com/Qiskit/rustworkx version 0.14.2 for evaluating Personalized PageRank on a CPU machine. For evaluation on a GPU machine, we used cuGraph  [^10] https://github.com/rapidsai/cugraph version 24.4.0. We used vllm  [^11] https://github.com/vllm-project/vllm version 0.4.2 for inference to generate summaries.